OpenStudy (theedwardsfamily):

*********Medal & Fan********* What is the solution to the equation? 1-z/5=9 A. z = 50 B. z = 44 C. z = 2 D. z = 40

OpenStudy (anonymous):

z=-40

OpenStudy (theedwardsfamily):

@Albany_Goon @KamiBug @Here_to_Help15 @sleepyjess @sanaybrown

OpenStudy (theedwardsfamily):

@pooja195

OpenStudy (alexandervonhumboldt2):

wait!!!!!!!!!1 forst 9-1=-z/5 -8=z/5 z=-40

OpenStudy (theedwardsfamily):

thnx guys!!!!!!!!!!!!!!! ;)